imdog wrote:
But Regicide receive 3 consecutive PK (though it is a bug). so the chance is 1/36 x 1/36 x 1/36 = 0.00000214. It only happened twice in a million game! (correct if I am wrong)
As you wish:
(1/36) x (1/36) x (1/36) is 0.0000214. twenty per million |